### v1

# On the Utility of Accounting for Human Beliefs about AI Behavior in Human-AI Collaboration

Abstract
To enable effective human-AI collaboration, merely optimizing AI performance while ignoring humans is not sufficient. Recent research has demonstrated that designing AI agents to account for human behavior leads to improved performance in human-AI collaboration. However, a limitation of most existing approaches is their assumption that human behavior is static, irrespective of AI behavior. In reality, humans may adjust their action plans based on their observations of AI behavior. In this paper, we address this limitation by enabling a collaborative AI agent to consider the beliefs of its human partner, i.e., what the human partner thinks the AI agent is doing, and design its action plan to facilitate easier collaboration with its human partner. Specifically, we developed a model of human beliefs that accounts for how humans reason about the behavior of their AI partners. Based on this belief model, we then developed an AI agent that considers both human behavior and human beliefs in devising its strategy for working with humans. Through extensive real-world human-subject experiments, we demonstrated that our belief model more accurately predicts humans’ beliefs about AI behavior. Moreover, we showed that our design of AI agents that accounts for human beliefs enhances performance in human-AI collaboration.
1 Introduction
The potential for human-AI collaboration is immense and spans various domains. In healthcare, AI systems can identify diagnoses that might be overlooked by human professionals (Cheng et al., 2016; Mobadersany et al., 2018). In industrial manufacturing, robots work alongside human workers to enhance efficiency and safety (Bauer et al., 2008; Sherwani et al., 2020). In workflow productivity, virtual assistants can generate drafts for humans to refine and finalize (Wilson and Daugherty, 2018). However, despite significant improvements in AI performance over the past decade, designing AI agents to optimize the overall performance of human-AI collaboration remains a challenge.
In particular, optimizing the AI system in isolation is not sufficient to enhance the performance of human-AI collaboration. Both the human and the AI agent need to coordinate by inferring the goals and intentions of their counterpart and taking complementary actions. For example, in AI-assisted decision-making, researchers have demonstrated that instead of optimizing AI by itself, training AI to focus on improving in areas where humans typically struggle can significantly enhance the performance of human-AI teams (Bansal et al., 2021; Wilder et al., 2021). In human-AI collaboration, Carroll et al. (2019) show that incorporating models of human behavior into the training of the AI agent leads to higher collaborative performance compared to training the agent to play with themselves through self-play (Silver et al., 2018). These studies highlight the importance of integrating human behavior into the design of collaborative AI. However, a key limitation of these research efforts is that they mostly assume that human behavior remains static, irrespective of the actions and behavior of the AI counterpart. In practice, humans may modify their behavior in response to their beliefs about what AI agents intend to do based on their observations of AI behavior.
Figure 1: Example task.
In this work, we argue that, in addition to incorporating human behavior, designing an AI agent that account for humans’ beliefs about AI behavior could significantly improve their collaborative performance. Consider a simple illustrative example in Figure 1, where the human and AI agents need to go to both goals $\{G1,G2\}$ to complete the task. Assume that the human prefers $G2$ over $G1$ because it requires fewer steps. When designing a collaborative AI that accounts for human behavior only, the AI would choose to go to $G1$, leading to lower overall collaborative performance. However, if the human would adjust their behavior based on their belief about AI behavior (e.g., they will avoid going to the same goal that the AI is going towards), we could incorporate this knowledge into the design of the AI. For example, the AI can choose to go to $G2$, anticipating that the human will account for that and go to $G1$, leading to improved overall performance. This example underscores the potential benefits of incorporating human beliefs about AI behavior in AI design. Meanwhile, it also highlights the importance of developing accurate models of human beliefs and incorporating them into the AI design.
In this work, we aim to examine the benefits of incorporating human beliefs about AI behavior in the design of AI to improve human-AI collaboration. Our work focuses on the setting where the human and AI agents need to coordinate their goals (i.e., they need to complete different subtasks in order to accomplish the overall task). More formally, we formulate the human-AI collaboration environment as a multi-player goal-oriented Markov decision process (MDP). We begin our investigation by developing models of human beliefs regarding AI behavior. This modeling effort extends the level-$k$ reasoning framework (Stahl II and Wilson, 1994) to account for suboptimal human behavior. Specifically, we first develop a behavioral level-0 model that assumes agents take actions without considering the behavior of other agents. We then enhance the belief model by introducing a behavioral level-1 model, which assumes humans interpret the behavior of another player as if the other agent adheres to the behavioral level-0 model. With the models of human behavior and beliefs established, we proceed to develop collaborative AI agents that incorporate different assumptions about human models.
To examine our proposed approaches, we conduct extensive human-subject experiments on two environments. To develop and assess the model of human behavior (i.e., the behavioral level-00 model), we first train models of human behavior using behavioral cloning on real-world human behavior and evaluate their performance. For the development and assessment of the model of human beliefs (i.e., the behavioral level-$1$ model), we conduct experiments that present each participant with a trace of behavior by an agent and ask them to infer the agent’s goal, examining whether our model leads to accurate predictions of human inference on the agent’s goal. Moreover, based on the belief model, we also explore the possibility of developing AI policies such that humans can more easily infer the goals of AI from its actions. Finally, utilizing the developed models of human behavior and beliefs, we conduct a final set of human-subject experiments that pair each human participant with an AI agent in a two-player coordination game. We assess the human-AI collaborative performance for different designs of AI agents. Overall, our results demonstrate the effectiveness of our developed model of human behavior and beliefs. Furthermore, we show that AI agents accounting for models of human behavior and beliefs achieve better collaborative performance with humans, compared to AI agents that do not consider human beliefs or those that disregard both human behavior and beliefs.
Contributions. The main contributions of our work can be summarized as follows:
  • $\bullet$
    We develop models of human beliefs, inspired by the level-$k$ framework, that account for human suboptimality. Moreover, we have designed AI agents that can take explicable action plans, making it easier for humans to infer AI goals from observing AI actions.
  • $\bullet$
    We have designed collaborative AI agents, considering both models of human behavior and human beliefs about AI behavior, aiming to optimize the performance of human-AI collaboration.
  • $\bullet$
    Through extensive human-subject experiments with 1,690 participants, we demonstrate that our belief models are effective and can help us design AI policies that are explicable to humans. Moreover, our design of collaborative AI that incorporates human beliefs about AI behavior improves the collaborative performance when paired with real-world human participants. The results demonstrate both the importance of incorporating human beliefs about AI behavior in the design of collaborative AI and the effectiveness of our approach.
2 Related Work
Our work joins the recent growing research in human-AI collaborations (Carroll et al., 2019; Jaderberg et al., 2019; Bansal et al., 2021; Wilder et al., 2021). Bansal et al. (2021) demonstrate that optimizing AI in isolation may lead to suboptimal performance for human-AI collaboration. Wilder et al. (2021) show that training AI systems to complement humans, performing better in areas where humans struggle, leads to improved collaborative performance. Carroll et al. (2019) illustrate that incorporating a human model, learned from human data, into the training of AI results in enhanced performance when these AI systems work with real humans. Additionally, Jaderberg et al. (2019) use population-based reinforcement learning to improve the robustness of trained AI agents. Our work extends this line of work by proposing to incorporate not only human behavior but also human beliefs of AI behavior into the design of collaborative AI.
From a technical perspective, our work involves understanding and modeling humans in decision-making. There has been a significant amount of work in the literature on modeling human behavior. For example, Inverse Reinforcement Learning (IRL)(Ng et al., 2000; Abbeel and Ng, 2004; Ramachandran and Amir, 2007) aims to infer the reward functions in Markov Decision Processes (MDPs) through observing demonstrations of the optimal policy. If the demonstrator is a human being, the demonstrations could be noisy or contain behavioral biases. Studies (Evans et al., 2016; Shah et al., 2019; Hughes et al., 2020; Zhi-Xuan et al., 2020) have aimed to incorporate human behavioral biases in the inference process and infer both the rewards and biases simultaneously. Imitation Learning (Hussein et al., 2017) also aims at developing models that can mimic human behavior from demonstrations. Recently, there have also been an increasing amount of research efforts that incorporates human models in computational and machine learning frameworks Kleinberg and Oren (2014); Tang and Ho (2019, 2021); Masters et al. (2021a, b); Yu and Ho (2022); Yu et al. (2023); Feng et al. (2024).
From the perspective of modeling human beliefs over others’ behavior, this has been discussed in the literature of level-$k$ reasoning (Stahl II and Wilson, 1994; Gill and Prowse, 2016) in economics and theory of mind (Premack and Woodruff, 1978; Westby and Riedl, 2023). As a few examples, Yu et al. (2006) have found that real human behavior is close to level-$1$ and level-$2$ reasoning models in cooperative games. Albrecht and Stone (2018) survey works on autonomous agents modeling the beliefs and intentions of other agents. They distinguish methods for modeling stationary or changing agent behaviors. The belief model includes theory of mind, recursive reasoning models, plan recognition, partially observable Markov decision processes (POMDPs), and others. Baker et al. (2011) propose a Bayesian theory of mind to describe human modeling of joint belief of state in a partially observable MDP and conduct human experiments in (Baker et al., 2009, 2017). A similar human modeling approach is adopted by Wu et al. (2021) and applied to the Overcooked experiment. Their idea is to split delivering a dish into several subtasks, such as picking up an ingredient or putting an ingredient into a pot, and the player is inferring others’ subtasks and selecting their own corresponding subtask before taking actions.
3 Problem Formulation and Methods
In this section, we first formulate the human-AI collaboration framework as a multi-player goal-oriented Markov decision process. We then outline our methods for modeling human behavior and beliefs about AI behavior. Finally, we present our approaches for integrating human behavior and beliefs into the development of collaborative AI agents.
3.1 Decision-Making Environment
We formulate the human-AI cooperative decision-making environment as a multi-agent Markov decision process (MDP), represented by $W=\langle S,\alpha,{A_{i}|_{i\in[\alpha]}},R,P\rangle$. $S$ is a finite set of states; $\alpha$ is a finite set of players; $A_{i}$ is the action set available to player $i$. $P:S\times A_{1}\times\dots\times A_{|\alpha|}\times S\rightarrow[0,1]$ is the transition function that determines the next state given all players’ joint actions. $R:S\times A_{1}\times\dots\times A_{|\alpha|}\rightarrow\mathbb{R}$ is the joint reward function assigned to all players given their joint actions. Note that in this work we consider the cooperative setting. Therefore, our formulation only incorporates a single reward function $R$ for all players, though this can be easily extended.
While our formulation could address cases with multiple human and AI players, in this work, we focus on a two-player cooperative game, i.e., $\alpha=\{1,2\}$, where one player is a human and the other is the AI. During the decision-making process, neither the human nor the AI knows the other player’s next action or future plans, and they cannot communicate directly. However, they can observe each other’s past actions, enabling them to infer about the other player and modify their own actions accordingly.
Goal-oriented MDP. In this work, we focus on the setting with goal-oriented MDP, where there is a set of goals $G=\{g_{1},\dots,g_{k}\}\subseteq S$, which are subset of states that are terminal states, i.e., $P(g|g,a)=1~{}\forall g\in G,a\in A$. Moreover, the decision-making agent only receives rewards when arriving at one of the goal states.
3.2 Modeling Human Behavior and Human Beliefs about AI Behavior
Our models are motivated by the level-$k$ framework (Stahl II and Wilson, 1994) in economics. In particular, we start by considering humans as level-00 agents that do not account for others’ behavior in the environment. Differing from the literature, we address the natural situation that level-00 agents do not behave optimally, and we call this model behavioral level-00 agents. To account for human beliefs, [^1] In this work, human beliefs refer to humans’ belief about the goal of the AI agent. we model humans as an extension of the level-$1$ agents, that assume other agents are behavioral level-00 agents [^2] Our model can iteratively progress to higher level-$k$ agents. However, we focus on the case with $k\leq 1$. and update their beliefs in a Bayesian manner based on the observations of others’ behavior.
Modeling human behavior. We first model human behavior under the assumption that humans do not consider other players in the environment (or that they consider other players as a part of the environment without strategically responding). In this case, a human behavior model can be represented as $H:W\rightarrow\Pi$, mapping a given environment $w\in W$ to a policy $\pi=H(w)$. We give two examples of human behavior models utilized in our work below.
  • $\bullet$
    Standard model. First consider the standard human behavior model in MDPs, in which the goal of the human is to maximize the expected cumulative reward, and their policy only depends on the current state. The model can be represented by $\pi(a|s)$, indicating the probability of choosing action $a$ at state $s$. For the standard model that assumes decision optimality, humans choose actions maximizing the Q-function, where $Q(s,a)$ indicates the expected cumulative reward if the player takes action $a$ in state $s$ and follows policy $\pi$. $Q(s,a)$ could be calculated by standard reinforcement learning techniques such as value iteration or Q-learning.
  • $\bullet$
    Behavioral level-00 model. We also consider the case that we can learn human behavioral models from their historical behavioral traces through behavioral cloning (Pomerleau, 1988; Torabi et al., 2018). Behavioral cloning is one of the imitation learning approaches, which learns a policy from human demonstration by building a map from states to actions with supervised learning methods Bain and Sammut (1995). We build a fully connected neural network, where the input is the state encoding, and output is the probability over action space, and train the model with standard gradient descent method with cross-entropy loss.
Modeling human beliefs. We now describe how we develop models for human beliefs. As summarized earlier, we consider cases where human decision-makers assume that other agents in the environments are behavioral level-00 agents. In our discussion, we also describe this belief model as behavioral level-$1$ agents.
[…]
$B(g|(s,a)_{1:t})\propto\lambda(g)Pr((s,a)_{1:t}|g)=\lambda(g)\prod_{1:t}\pi(a_{i}|s_{i},g)P(s_{i}|s_{i-1},a_{i-1})$ (1)
3.3 Designing Collaborative AI Agents
To illustrate the effectiveness of incorporating models of human beliefs, we train different AI agents that work with humans in human-AI collaboration via simulations and real-world human subject experiments.
Training methodology. The main idea of our training method is through self-play. We incorporate the models of humans and have AI teammates play with the agents specified by the human models through simulated plays. We use proximal policy optimization (PPO) to train collaborative AI agents.
Collaborative AI agents. Using the methodology above, we have designed several AI agents based on different assumptions of the behavior of human counterparts.
  • $\bullet$
    Assuming humans are optimal. We first train an AI agent that learns to collaborate with itself through self-play. This is equivalent to assuming the human counterpart is acting optimally.
  • $\bullet$
    Assuming humans are behavioral level-00 agents. We next train an AI agent that assumes the human is a behavioral level-00 agent, using behavioral cloning to train the behavioral model.
  • $\bullet$
    Incorporating models of human behavior and beliefs. Finally we also train an AI agent that incorporates both the models of human behavior and beliefs into the design of AI agents.
The details of the setup and implementations are included in the appendix.
4 Experiments
To evaluate our approaches, we have conducted three sets of experiments, consisting of both simulations and human-subject experiments. For our human-subject experiments, we have recruited in total $1,690$ participants from Amazon Mechanical Turk (MTurk) for multiple sets of experiments. The experiments are approved by the IRB of our institution. Workers were paid \$1 base payments with the potential for bonus payments in some experiments. The average hourly rate is approximately \$$\$14$ across all our experiments.
4.1 Experiment Environment: Grid World with Two Players and Multiple Goals
Our experiments are conducted in a grid world environment with two players and multiple goals. We have conducted two variations of the experiments. In the first variation, we designed the environments such that the two players are not playing in the same grid world. However, each of them has access to the full information of the environment and the actions of the other player. This variation abstracts away the interdependency of agent actions, meaning the agents’ actions do not influence each other directly. This allows us to focus on how humans and AI reason about each other’s goals and intentions. Note that we obtained qualitatively similar results in this first variation and the second, more complicated variation of environments described next. For simplicity, we have included the setup and results of the first variation in the appendix. Below, we describe the experiment environment and results of our second variation.
More specifically, as shown in Figure 2, in our second variation of the environment, the grid world is 8 by 8 in size, containing the positions of both players and four possible goals. The players can choose to move {Up, Down, Right, Left} or stay in the current grid. They can see each other’s positions and take actions simultaneously. Each player can navigate to two of the four goals. In particular, the human player can only reach one of the two goals colored blue, and the AI player can only reach one of the goals colored green. When both agents reach the same type of goal (both reaching a "star" or both reaching a "triangle"), they earn positive points. However, they will not earn points if they reach different goals or if they collide into each other (move into the same position). We set the maximum number of actions to be 20.
(a) The interface for experiment 1 & 3.
(b) The interface for experiment 2.
Figure 2: The interface for our human-subject experiment. In experiment 1, each participant is playing with an AI agent. The participants in experiment 1 are told what their goal is and only need to focus on reaching the goal without colliding with the AI agent. In experiment 2, each participant is provided traces of the behavior by other agents and is asked to infer which goal one agent is trying to reach. Finally, in experiment 3, the participants are not told which goal to reach, and they need to make decisions based on their beliefs over AI behavior. In experiments 1 and 3, the participants can only receive bonus rewards by reaching the same type of goals (star or triangle) as the AI agent.
In the first experiment, we recruit real-world humans to play against another pre-defined AI agent. We then examine the effectiveness of utilizing behavioral cloning as a model to mimic human behavior. Note that using behavioral cloning to model human behavior is extensively adopted in the literature, and the purpose of the first experiment serves as the foundation to develop our human belief models and the design of collaborative AI agents.
To evaluate our proposed approaches in modeling human beliefs and designing cooperative AI, we have designed and conducted additional two sets of experiments. In the second set of experiments, we provide participants with different traces of actions from other agents, and ask participants to infer the goal of the agents. This experiment helps us evaluate whether our belief model leads to better predictions of human beliefs over others’ behavior. Afterwards, in the third experiment, we conduct experiments similar to the first experiment. We team up AI agents and human players to examine the team performance of our design of collaborative AI when working with real-world human participants. But unlike the first experiment, we don’t reveal any information about goals the players are trying to reach.
4.2 Experiment 1: Training and Evaluating Human Behavior Models
In our first experiment, our goal is to examine the effectiveness of utilizing behavioral cloning to model human behavior. Note that his approach has been proposed and examined in various domains Carroll et al. (2019). Our main purpose to conduct this experiment is to ensure whether this method works well in our setting, as our proposed approaches in developing human belief models and collaborative AI are built on top of models of human behavior.
Experiment setup. We recruited 190 workers from Amazon Mechanical Turk. Each recruited worker was asked to play 30 navigation games within a grid world with a pre-defined AI model, as shown in Figure 2a. To collect diverse data, we designed three AI models to play with humans: random AI (selecting a random goal and taking the shortest path), self-play AI (trained using self-play), and optimal AI (optimized via joint optimization).
Note that in this experiment, our goal is to develop a model of human behavior that outputs a sequence of actions given a goal. Therefore, in experiment 1, we provided additional instructions to participants about which goal they should reach in each round, and the AI models controlled the other player to move towards the same type of goals. In our later experiment 3, we did not provide these instructions, and humans needed to infer the goal of the AI agent.
Evaluation of data-driven behavior models. We divided the collected data of human actions into three sets: training, validation, and testing. The training set comprised data from 152 workers, including approximately 136,000 instances of user decisions, while the validation and testing sets each contained data from 19 workers, amounting to around 17,000 instances of user decisions each. We employed a 4-layer Multilayer Perceptron (MLP) model, where the input is the current environment layout, and the output predicts the next human action. We fine-tuned the hyperparameters, such as learning rate, hidden layer size, and L2 penalty, based on validation errors.
Our results suggest that human behavior is noisy and is significantly away from being optimal. Even when human participants are provided with suggested goals during data collection, there is only a $55.0\%$ chance that both players will reach the same type of goals across all treatments. Besides, there is a notable chance of two players colliding (about $15.8\%$) or ending in different types of goals (about $25.6\%$). We compared the performance of our learned model to a model predicated on optimal agent behavior, defined as taking the shortest path to the goal. The training, validation, and test accuracies of both models are presented in Table 1. These results clearly reveal that human behavior deviates significantly from the assumed optimality. This deviation highlights the importance of incorporating a realistic model of human behavior into human-AI cooperation frameworks.
Table 1: The prediction accuracy for human behavior assuming optimal behavior and using data-driven model in Experiment 1.
Training Accuracy Validation Accuracy Testing Accuracy
Assuming Optimal Behavior 0.4498 0.4327 0.4459
Data-Driven Model 0.8547 0.7831 0.7899
4.3 Experiment 2: Evaluating Human Belief Models
We next examine our model of human beliefs. Specifically, we investigate whether we can design AI behavior such that it is easier for humans to infer the goal of the AI agent.
In our experiments with a simplified environment, as included in the appendix, we directly examined how accurately our belief model can infer human beliefs by comparing the predictions of our models with the direct solicitation of users’ inferences about others’ behavior. The results suggest that our model is more accurate in predicting human beliefs compared to baselines. However, human beliefs are generally very noisy, and if we randomly draw a behavior trace from another agent and ask humans to predict the goal, human prediction accuracy is close to random guessing. Therefore, we next shift our focus to whether AI agents can design their action plans to make it easier for humans to infer the goal. In this more complicated environment, we have directly examined this follow-up question of designing AI policies that are explicable to humans.
Experiment setup. The experiment setup is presented in Figure 2b. In addition to player positions and goal positions, we also display behavioral traces, which are sequences of actions taken by previous players. For each participant, we show them the traces of two players and ask the participant to infer which one of the two goals those agents are trying to reach.
We compare two belief models. The first one is the standard level-1 model: humans assume the other agent is a level-0 agent that takes the optimal decision (i.e., taking the shortest path towards the goal). The second one is our proposed behavioral level-1 model: humans assume the other agent is also a suboptimal decision-maker, taking actions in the same way as themselves. In our implementation, we leverage the data-driven behavioral model derived from Experiment 1 as the user behavioral model. More concretely, we construct a belief model using Bayesian inference, as described in Section 3.2, using the corresponding human model of behavior.
Explicable AI policy. As mentioned earlier, humans generally struggle to infer the goals of other agents from behavioral traces randomly drawn from history (see the appendix for detailed results). Therefore, in this experiment, instead of displaying a randomly drawn historical behavioral trace, we display the behavioral trace of an AI agent that aims to make its behavior explicable. In particular, based on the developed belief models, we train AI agents to not only achieve their goals but also to maximize the likelihood that humans can accurately infer these goals from their behavior. Our implementation utilizes self-play and awards an additional bonus when the goal inference aligns with the predictions of a human belief model, in addition to the rewards for goal achievement. The bonus is proportional to the log likelihood of belief model inference $log(\pi(a|s,g))$. By conducting experiments with this explicable AI policy, we simultaneously evaluate whether our belief models are accurate and whether our design of explicable AI policy is effective.
Figure 3: The simulation results in Experiment 2.
Simulations. We design the AI agents to incorporate the belief models of standard level-1 agents and behavioral level-1 agents, enabling them to adopt policies that simplify the task of goal inference for humans. For each environment with a given goal, we generated the behavioral traces of both AI policies. The goal is to examine whether the explicable AI with our belief model makes it easier for humans to infer the goal of the AI. We first conducted simulations assuming humans are behavioral level-1 agent. The results are shown in Figure 3. The explicable AI equipped with the belief model of behavioral level-1 agents leads to behavioral traces that are easier to infer. Note that this result is not surprising is not surprising since we assume humans infer the goal following the beliefs of behavioral level-1 agent. The results provide evidence that our design of explicable AI is effective.
Figure 4: The results of human-subject experiments in Experiment 2.
Human-subject experiments. We next conducted human-subject experiments to assess whether our approach results in actions that make it easier for real humans to infer the goal. We recruited 200 workers from Amazon Mechanical Turk, randomly assigning them to one of two treatments. Each participant was tasked with identifying the goals of the player in 30 different scenarios. The length of actions is drawn from the range 4 to 8. To incentivize effort, participants were awarded a \$$\$0.03$ bonus for each correctly identified goal. The results, as shown in Figure 4, demonstrate that humans are better at inferring the goal of the explicable AI agent, that assumes humans are behavioral level-1 agents. The difference is statistical significant with $p>0.0001$. This comparison demonstrates that when coupled with the more accurate human belief models, we can indeed design AI policy to be explicable, i.e., making it easier for humans to infer AI goals from observing AI actions.
4.4 Experiment 3: Evaluating the Design of Collaborative AI Agents
We now examine our design of collaborative AI agents. As described in Section 3.3, we train collaborative AI agents using three different corresponding human models and utilize both simulations and human-subject experiments to examine the effectiveness of our design.
Experiment setup. The setup of our Experiment 3 is the same as Experiment 1, shown in Figure 2a. But human participants will not receive any hint about which goal their teammate or themselves are suggested to reach.
Simulations. We first run simulations to evaluate the performance of human-AI teams with different design of collaborative AI. The evaluation is based on $10,000$ randomly generated environments. However, we further filter out cases where the distance between the two goals for the same player is smaller than 3 to encourage models to adjust their behavior based on the inference of their teammate’s actions. We examine the performance of three designs of collaborative AI using the methodology described in Section 3.3:
  • $\bullet$
    Self-Play is the AI that is trained assuming they are playing with itself.
  • $\bullet$
    Behavior-AI is the AI that assumes the human partner is the behavioral level-00 agent.
  • $\bullet$
    Belief-AI is the AI that assumes the human partner is the behavioral level-$1$ agent.
We partner the collaborative AI agents with the three simulated human agents and measure the collaborative performance. The simulation results, shown in Table 2, highlight the importance of selecting appropriate human models for training collaborative AI. When an AI agent is paired with the human model that is used to train the AI, the collaborative performance is better compared to pairing with other models.
Table 2: Simulation results of collaborative performance over 10k testing cases in experiment 3. Column players are different AI agents, and row players are different simulated human models.
Human Model AI Agent
Self-play Behavior-AI Belief-AI
Self-play AI 0.6245 0.5250 0.6177
Behavior Model 0.4902 0.6334 0.5755
Behavior&Belief 0.6411 0.6574 0.7675
Figure 5: The average human-AI performance in our human-subject experiment in Experiment 3.
Human-subject experiments. To examine the performance of our collaborative AI design when pairing with real humans, we recruited 200 workers and divided them into three treatment groups, each interacting with a different design of AI. Participants were tasked with playing 30 games, preceded by three tutorial games designed to familiarize them with the gameplay. Participants can receive a \$0.05 bonus payment for reaching the same goal with the AI agent for each of the 30 tasks.
Figure 5 presents the average collaborative reward, highlighting that AI trained with models of human behavior and beliefs achieved the highest collaborative performance when working with real humans. This outcome not only validates our design of AI but also suggests that real humans’ actions align with the behavior and belief models in this environment. Statistical analysis shows significant differences in the performance of AI models paired with humans, with $p<0.0001$ for comparisons between self-play and AI accounting for human behaviors, and a $p$ value of $0.0014$ for comparisons between AI accounting for human behaviors versus accounting for both behavior and belief. These results demonstrate that incorporating human beliefs into the deign of AI agents enhances collaborative performance when working with real humans.
5 Conclusion
Our work explores the impact of incorporating human behavior and beliefs into the design of collaborative AI, aiming to improve collaboration between humans and AI. By developing models that account for human beliefs regarding AI actions and integrating these models into AI design, we have observed improved performance in human-AI collaboration. Our approach, validated through simulations and human-subject experiments, demonstrates that AI agents designed with an understanding of human behavior and beliefs can be more effective in working with humans. This suggests a potential path forward for creating AI systems that are better aligned with human partners, making collaborative tasks more efficient and intuitive.
Generalization and limitations. We have demonstrated the effectiveness of our approach through human-subject experiments in commonly used grid world environments with varying levels of complexity. However, as with most human-subject studies, our findings are limited to the chosen environments. Moreover, our approach leverages historical data to construct models of human behavior and beliefs. This approach implicitly assumes that both human behavior and beliefs remain unchanged over time. Therefore, it is important to improve the development of human models and examine their generalizability to other environments.
Societal impacts. Our approach highlights the importance of incorporating not only human behavior but also human beliefs about AI behavior in the design of AI for human-AI collaboration. We believe this will become increasingly important as AI capabilities grow. However, while our research focuses on enhancing human-AI collaboration, there is a potential for the same methods to be used negatively, such as designing AI that intentionally sabotages human utility. One approach to mitigate potential negative impacts is to increase the transparency of AI models, enabling humans to develop appropriate reliance when working with AI agents.
Acknowledgements
This work is supported in part by J.P. Morgan Faculty Research Award and a Global Incubator Seed Grant from McDonnell International Scholars Academy.
Appendix A Additional Experiments
Besides experiments in Section 4, we conduct additional experiments in a simple grid world environment with multiple goals to illustrate our study of humans’ beliefs about AI behavior.
A.1 Experiment Environment: Grid World with Multiple Goals
We use grid worlds of size of $6\times 6$ in both simulations and human experiments. Similar to the environment setup in Section 4.1, the grid world contains a start position, two goal positions, and some blocked positions that the player cannot enter. The player needs to move from the start position towards one of the goal positions. The player can choose to move {Up, Down, Right, Left}. The player will get a positive reward upon reaching the goal, and we set the maximum number of actions to be 20.
We first recruit participants to engage in a single-player game and record their behavior. We then build human behavior model using behavioral cloning. To evaluate our proposed approaches in modeling human beliefs and designing cooperative AI, we conducted additional two sets of experiments. In the second set of experiments, we provide participants with different traces of actions from other agents, again in a single-player game, and ask participants to infer the goal of the agents. This experiment helps us evaluate whether our belief model leads to better predictions of human beliefs over others’ behavior. Afterwards, we conduct a two-player game in the third experiment, where humans are paired with different AI agents to examine the team performance of our design of collaborative AI.
The interfaces of our experiments can be seen in Figure 6a, 6b, and 6c. The detailed descriptions of the experiment and interfaces are included in Appendix C.
(a) Experiment 4.
(b) Experiment 5.
(c) Experiment 6.
Figure 6: Human-subject experiment interfaces. In Experiment 4, each participant is asked to control the player to move to the goal (red star). In Experiment 5, each participant is provided a trace of the behavior by another agent, and is asked to infer which goal the agent is trying to reach. In Experiment 6, each participant is playing with an AI agent in separate environments. The participants only receive bonus rewards by reaching the same goal (red star or green triangle) as the AI agent.
A.2 Experiment 4: Training and Evaluating Human Behavior Models in Single-Player MDP
Similar to Experiment 1, we recruited 200 workers from Amazon Mechanical Turk. Each recruited worker was asked to play 15 navigation games within a grid world, as shown in Figure 6a. Our goal is to leverage the collected data to create a data-driven model of human behavior. We divided the collected data of human actions into three sets: training, validation, and testing. The training set comprised data from 160 workers, including approximately 70,000 instances of user decisions, while the validation and testing sets each contained data from 20 workers, amounting to around 8,800 instances of user decisions each. The model and parameter tuning were similar to what we did in Experiment 1.
Evaluation of data-driven behavior models. The training, validation, and test accuracies of assuming optimal behavior and behavioral models are presented in Table 3. Results show that data-driven model could predict human behavior more accurately than assuming humans are optimal.
Table 3: The prediction accuracy for human behavior for different human models in Experiment 4.
Training Accuracy Validation Accuracy Testing Accuracy
Assuming Optimal Behavior 0.7266 0.6964 0.7131
Data-Driven Model 0.9189 0.8136 0.8422
A.3 Experiment 5: Evaluating Human Belief Models
We developed human belief models (behavioral level-1) similar to previous experiments. We examine whether belief inference using behavioral model aligns with real humans, and whether we can design AI behavior such that it is easier for humans to infer the goal of the AI agent.
Experiment setup. The experiment setup is presented in Figure 6b. The grid world contains a starting position and two goal positions. For each participant, we show them a trace of behavior from another agent and ask the participant to infer which of the two goal the agent is trying to reach.
Experiment 5.1: Examining the belief models. We recruited 200 workers from Amazon Mechanical Turk to compare standard level-$1$ and behavioral level-$1$ model. Each worker was asked to infer the goal for 25 behavioral traces. We compared the performance of two belief models and the worker accuracy, as shown in Table 4. As we can see from the table, humans are generally poor at inferring the goal of other agents: their accuracy only reaches $59.37\%$ in inferring the goal of the other agent. Moreover, the behavioral level-$1$ model, which accounts for the human behavior model in the belief model, captures human beliefs better than the standard level-$1$ model, which assumes human behavior is optimal.
Table 4: Performance comparison between Bayesian inference framework assuming standard model and human behavior model.
Consistency to Human Predictions Cross Entropy Loss
Standard level-$1$ 0.4977 0.9284
Behavioral level-$1$ 0.5764 0.7506
True goals 0.5937 0.6631
Figure 7: Human evaluations of belief inference accuracy regrading AI goals.
Experiment 5.2: Developing explicable AI policy. As demonstrated in Experiment 5.1, humans generally struggle to infer the goals of other agents based on others’ behavior. To design AI agents which makes inference of their goals easier by humans, we train AI agents to maximize the likelihood that humans can accurately infer these goals from their behavior.
We recruited 400 workers from Amazon Mechanical Turk, randomly assigning them to one of two treatments, to assess the behavior of AI agents. Each participant was tasked with identifying the goals of the player in 30 different scenarios. Participants were awarded a \$$\$0.03$ bonus for each correctly identified goal. Figure 7 displays the results of human evaluation. The findings indicate that humans achieve higher accuracy in inferring the correct goals of AI agents when employing Behavioral level-$1$ model.
A.4 Experiment 6: Evaluating the Design of Collaborative AI Agents
Utilizing developed models of human behavior and beliefs, we follow the same methodology in Experiment 3: train different collaborative AI, pair them with different human models, and examine the collaborative performance in simulations and human subject experiments (via recruiting 300 workers).
The setup of our Experiment 6 is shown in Figure 6c. The goal for the human-AI team is for both agents to reach the same goal in their own environments (both reaching “red star” or both reaching “green triangle”) within a time limit. The team will not get points if they reach different goals or one of the players fails to reach any goal.
Simulations results are shown in Table 8. Simulations indicate that collaborative performance is highest when an AI agent is paired with the human model used to train it. Figure 8 presents the average collaborative reward in human-subject experiments. Statistical analysis revealed significant differences in the performance of AI models when paired with human participants, with $p$-values of $0.0166$ for comparisons between the treatments self-play and Behavior-AI, and $p<0.0001$ for Behavior-AI versus Belief-AI. These results demonstrate that incorporating human beliefs into the deign of AI agents enhances collaborative performance when working with real humans.
[…]
tableSimulation results of human-AI collaborative rewards in Experiment 6. Columns players are different AI agents, and row players are different simulated human models. Human Model AI Agent Self-play Behavior-AI Belief-AI Self-play AI 0.7828 0.4780 0.6164 Behavior 0.5926 0.7584 0.4552 Behavior&Belief 0.6919 0.7268 0.7813
Figure 8: Average collaborative reward in human-subject experiment 6.
Appendix B Implementation Details
All our experiments are conducted on a cluster of 40 CPU cores (Intel Xeon Gold 6148 CPU @@@@ 2.40GHz), 2 GPUs (NVIDIA Tesla V100 SXM2 32GB), and a maximal memory of 80GB. For completeness, we provide the details of the implementation and the tuning of hyperparameters below.
State encoding of grid world. For a grid world of size $N\times N$ and contains $k$ different objects, we use $k$ channels to encode the environment state. Each channel is a binary matrix of size $N\times N$, indicating whether it contains specific objects. For example, in a single-player environment with one goal (environment in Experiment 4), we encode space, blocks, the start position, and goal positions as four channels. Thus the input state is of size $4\times 6\times 6$ before flattening. The state encoding is used as input of our behavioral cloning models and AI agents.
Behavioral cloning for grid world of size $6\times 6$. We construct a neural network with four fully connected layers to model human behavior. The input is the state encoding, represented as a matrix of dimensions $4\times 6\times 6$, and the output is the probability of taking various actions. The network parameters are initialized using the Glorot uniform initializer. We employ the Adam optimizer and cross-entropy loss to optimize the neural network parameters. The batch size is set to $1024$, and the maximum number of training epochs is set to $1000$. Training is halted early if the validation loss begins to increase. The hyperparameters are adjusted based on the performance on the validation dataset. The number of nodes for each hidden layer is tuned within the range $\{64,128,256,512,1024\}$, the initial learning rate within the range $\{0.001,0.003,0.01,0.03,0.1\}$, and the L2 penalty within the range $\{0,0.01,0.03,0.1\}$. Each training process takes $10$ to $30$ minutes depends on hyperparameters. In our reported results, we choose hidden layer size of $512$, learning rate $0.01$ and L2 penalty $0.01$.
Behavioral cloning for grid world of size $8\times 8$. The neural network structure mirrors that used for the grid world of size $6\times 6$. The input is the state encoding, represented as a matrix with dimensions $8\times 8\times 8$, and the output is the probability of taking various actions. The initialization, optimizer, loss function, batch size and training process are consistent with those used in the grid world of size $6\times 6$. The maximum of training epochs is set to $3000$. The hyperparameters are adjusted based on the performance on the validation dataset. The number of nodes for each hidden layer is tuned within the range $\{1024,2048,4096\}$, the initial learning rate within the range $\{0.001,0.003,0.01,0.03,0.1,0.3\}$, and the L2 penalty within the range $\{0,0.01,0.1\}$. In our reported results, we choose hidden layer size of $2048$, learning rate $0.01$ and L2 penalty $0.01$.
AI agent training for grid world of size $6\times 6$. We utilize the Proximal Policy Optimization (PPO) and actor-critic framework for our AI model. Both the actor and critic networks employ four fully connected layers, maintained at the same size. The input is the state encoding of a two-player game, which includes two game states and the index vector (the input vector length is $290$). The critic network outputs a single value, while the actor network produces a distribution over the action space. The batch size is set to $256$, and the maximum number of training episodes is limited to $1000$. The discount factor is set to $0.99$. For PPO optimization, the clipping parameter $\epsilon$ is fixed at $0.2$, and the number of PPO epochs is $10$. The number of nodes for each hidden layer is tuned within the range $\{64,128,256,512,1024\}$, and the initial learning rate is adjusted within the range of $\{0.001,0.003,0.01,0.03,0.1\}$. Each training process takes about $2$ to $6$ hours depends on hyperparameters. In comparisons between the self-play AI and the AI trained to play with a human model, we use the same model size (hidden layer size of $256$) and the same initial learning rate of $0.01$.
AI agent training for grid world of size $8\times 8$. The model structure mirrors that of the grid world of size $6\times 6$. The input size is adjusted to accommodate the new state representation (the input vector length is $514$). The training batch size, discounting factor, maximum number of training epochs and clipping parameters remain unchanged. The number of nodes for each hidden layer is tuned within the range $\{256,512,1024,2048\}$, and the initial learning rate is adjusted within the range of $\{0.01,0.03,0.1,0.3,0.5\}$. In comparisons between the self-play AI and the AI trained to play with a human model, we consistently use the same model size (hidden layer size of $512$) and the same initial learning rate of $0.1$.
Appendix C Human Experiment Details
We include more details about human-subject experiments. The experiment interface is shown in Figure 9. We conduct four human-subject experiments, and recruited 1690 workers from Amazon Mechanical Turk in total. Table 5 lists the number of workers, the number of tasks per worker and the number of treatments in each experiment. Table 6 contains the demographic information of all the workers.
(a) Experiment 1 & 3 interface: humans play games with AI agents in grid worlds of size $8\times 8$.
(b) Experiment 2 interface: humans infer goals of other players in grid worlds of size $8\times 8$.
(c) Experiment 4 interface: humans play single-player navigation games in grid worlds of size $6\times 6$.
(d) Experiment 5 interface: humans infer goals of other players in grid worlds of size $6\times 6$.
(e) Experiment 6 interface: humans play games with AI agents in grid worlds of size $6\times 6$.
Figure 9: Human-subject experiment interfaces.
Table 5: Number of workers in each human-subject experiment.
Workers Tasks per Worker Treatments Base Payment Bonus per Task
Experiment 1 190 30 3 \$1.00 \$0.05
Experiment 2 200 30 2 \$1.00 \$0.03
Experiment 3 200 30 3 \$1.00 \$0.05
Experiment 4 200 15 1 \$1.00 0
Experiment 5.1 200 25 1 \$1.00 0
Experiment 5.2 400 30 2 \$1.00 \$0.03
Experiment 6 300 20 3 \$1.50 \$0.05
Table 6: Demographic information of all the participants in our experiments.
Group Category Number
Age 20 to 29 775
30 to 39 642
40 to 49 146
50 or older 127
Gender Female 547
Male 1118
Other 25
Race / Ethnicity Caucasian 1511
Black or African-American 82
American Indian/Alaskan Native 32
Asian or Asian-American 28
Spanish/Hispanic 7
Other 30
Education High school degree 60
Some college credit, no degree 42
Associate’s degree 39
Bachelor’s degree 1217
Graduate’s degree 293
Other 39