### v2

# On the Utility of Accounting for Human Beliefs about AI Intention in Human-AI Collaboration

Abstract.
To enable effective human-AI collaboration, merely optimizing AI performance without considering human factors is insufficient. Recent research has shown that designing AI agents that take human behavior into account leads to improved performance in human-AI collaboration. However, a limitation of most existing approaches is their assumption that human behavior remains static, regardless of the AI agent’s actions. In reality, humans may adjust their actions based on their beliefs about the AI’s intentions, specifically, the subtasks they perceive the AI to be attempting to complete based on its behavior. In this paper, we address this limitation by enabling a collaborative AI agent to consider its human partner’s beliefs about its intentions, i.e., what the human partner thinks the AI agent is trying to accomplish, and to design its action plan accordingly to facilitate more effective human-AI collaboration. Specifically, we developed a model of human beliefs that captures how humans interpret and reason about their AI partner’s intentions. Using this belief model, we created an AI agent that incorporates both human behavior and human beliefs when devising its strategy for interacting with humans. Through extensive real-world human-subject experiments, we demonstrate that our belief model more accurately captures human perceptions of AI intentions. Furthermore, we show that our AI agent, designed to account for human beliefs over its intentions, significantly enhances performance in human-AI collaboration.
Key words and phrases:
Human-AI Collaboration, Reinforcement Learning, Belief Inference
footnotetext: Preprint. Under review.
1. Introduction
The potential for human-AI collaboration is immense and spans various domains. In healthcare, AI systems can identify diagnoses that might be overlooked by human professionals (Cheng et al., 2016; Mobadersany et al., 2018). In industrial manufacturing, robots work alongside human workers to enhance efficiency and safety (Bauer et al., 2008; Sherwani et al., 2020). In workflow productivity, virtual assistants can generate drafts for humans to refine and finalize (Wilson and Daugherty, 2018). However, despite significant improvements in AI performance over the past decade, designing AI agents to optimize the overall performance of human-AI collaboration remains a challenge.
In particular, optimizing the AI system in isolation is often shown to be insufficient for enhancing the performance of human-AI collaboration. To optimize collaborative performance, AI agents need to account for their human counterparts when deciding on their strategies. For example, in AI-assisted decision-making, researchers have demonstrated that, instead of optimizing AI independently, training AI to focus on improving in areas where humans typically struggle can significantly enhance the performance of human-AI teams (Bansal et al., 2021; Wilder et al., 2021). In human-AI collaboration, Carroll et al. (2019) show that incorporating models of human behavior into the training of AI agents leads to higher collaborative performance compared to training the agent through self-play (Silver et al., 2018). These studies highlight the importance of integrating human behavior into the design of collaborative AI. However, a key limitation of these research efforts is that they mostly assume human behavior remains static, irrespective of the actions and behavior of the AI counterpart. In practice, humans may modify their behavior in response to their beliefs about the intentions of AI agents based on their observations of AI behavior.
In this work, we argue and aim to demonstrate that, in addition to incorporating human behavior, designing an AI agent that accounts for humans’ beliefs about AI intentions can significantly improve their collaborative performance. To illustrate this argument, consider a simple example task in Figure 1. In this task, both the human and AI need to collaborate to put out a fire in the center of the map. To successfully extinguish the fire, they need to use both a fire extinguisher and a hydrant, with the human and AI agent each retrieving one of them. Assuming they cannot directly communicate but can observe the direction the other is moving towards, if the AI agent believes the human will simply take the tool closest to them (the hydrant) and will not adjust their actions based on the AI’s behavior, the AI would choose to take the longer route to get the fire extinguisher. However, if the human would actually update their actions based on their inference of what the AI is trying to do from observing the AI’s actions, the more efficient strategy for the AI would be to move towards the hydrant and let the human grab the fire extinguisher. This example task demonstrates that, if humans would update their behavior based on their beliefs over AI intentions, we can design AI that account for human belief to improve overall human-AI performance.
(a) When the AI agent does not account for human beliefs about AI intentions and only considers human behavior, it would assume that the human would grab the closer hydrant. Therefore, the AI would go for the fire extinguisher, which would take 12 steps to put out the fire.
(b) When the AI agent accounts for human beliefs over AI intentions, it will try to reveal its intentions for humans so humans can take a complementary tool that would reach the overall better outcome. In this case, it takes 9 steps for the fire to be put out.
Figure 1. An example human-AI cooperation task that demonstrates the benefits for the AI to account for human beliefs about AI intentions.
In this work, we investigate and examine the benefits of incorporating human beliefs about AI intentions in the design of AI to improve human-AI collaboration. To operationalize the concept of AI intentions, our work focuses on a setting where human and AI agents need to coordinate their goals (e.g., they need to complete different subtasks) to accomplish the overall task. In this setting, the intention of the AI is defined as the goal or subtask the AI is attempting to achieve.
More formally, we formulate the human-AI collaboration environment as a multi-player goal-oriented Markov decision process (MDP). We begin our investigation by developing models of human beliefs regarding AI intentions. This modeling effort extends the level-$k$ reasoning framework (Stahl II and Wilson, 1994) to account for suboptimal human behavior. Specifically, we first develop a behavioral level-0 model that assumes agents take actions without considering the behavior of other agents. We then develop the belief model by introducing a behavioral level-1 model, which assumes humans interpret the behavior of another player as if the other agent take actions according to the behavioral level-0 model. With the models of human behavior and beliefs established, we proceed to develop collaborative AI agents that incorporate different assumptions about human models.
To examine our proposed approaches, we conduct extensive human-subject experiments. To develop and assess the model of human behavior (i.e., the behavioral level-00 model), we first train models of human behavior using behavioral cloning on real-world human data and evaluate their performance. For the development and assessment of the model of human beliefs about AI intentions (i.e., the behavioral level-$1$ model), we conduct experiments that present each participant with a trace of behavior by an agent and ask them to infer the agent’s goal, examining whether our model accurately predicts human inference of the agent’s intention. Moreover, based on the belief model, we explore the possibility of developing AI policies that make it easier for humans to infer the intentions of AI from its actions. Finally, utilizing the developed models of human behavior and beliefs about AI intentions, we conduct a final set of human-subject experiments in which each human participant pairs with an AI agent in a two-player coordination game. We assess the human-AI collaborative performance for different designs of AI agents. Overall, our results demonstrate the effectiveness of our developed models of human behavior and beliefs about AI intentions. Furthermore, we show that AI agents accounting for models of human behavior and beliefs about AI intentions achieve better collaborative performance when working with humans compared to AI agents that either do not consider human beliefs or disregard both human behavior and beliefs.
Contributions. In this work, we demonstrate that incorporating human beliefs about AI intentions in the design of collaborative AI can significantly improve human-AI collaborative performance. More concretely, the main contributions of our work are summarized as follows:
  • $\bullet$
    We develop models of human beliefs about AI intentions, inspired by the level-$k$ framework, and extend this framework to account for human suboptimality. In addition, we design AI agents capable of executing explicable action plans, making it easier for humans to infer AI intentions based on observed AI actions.
  • $\bullet$
    We design collaborative AI agents that incorporate both models of human behavior and human beliefs about AI intentions, aiming to optimize the performance of human-AI collaboration.
  • $\bullet$
    Through extensive human-subject experiments, we demonstrate that our belief models are effective and can guide the design of AI policies that are explicable to humans. Moreover, our design of collaborative AI that integrates human beliefs about AI intentions significantly improves collaborative performance when paired with real-world human participants. These results underscore both the importance of incorporating human beliefs about AI intentions in the design of collaborative AI and the effectiveness of our approach.
2. Related Work
Our work contributes to the growing body of research on human-AI collaboration (Carroll et al., 2019; Jaderberg et al., 2019; Bansal et al., 2021; Wilder et al., 2021). For example, Bansal et al. (2021) demonstrate that optimizing AI in isolation may lead to suboptimal performance in human-AI collaboration, compared to training AI with consideration for human performance. Wilder et al. (2021) show that training AI systems to complement humans by performing better in areas where humans struggle results in improved collaborative performance. Carroll et al. (2019) illustrate that incorporating a human model, learned from human data, into the training of AI enhances performance when these AI systems collaborate with real humans. Our work extends this line of research by not only accounting for human behavior when designing AI but also incorporating human beliefs about AI intentions.
While these approaches require knowledge or data of humans, there have been zero-shot coordination techniques being proposed to train cooperative AI agents when human data are not available. Jaderberg et al. (2019) use population-based reinforcement learning to improve the robustness of trained AI agents. Strouse et al. (2021) develop a variety of self-play agents by varying random seeds of neural network initialization and various checkpoints, and train a best response model tailored to the agent population, discovering that humans prefer to partner with this model. Zhao et al. (2023) utilize the advances of maximum entropy reinforcement learning to encourage the diversity of agent populations and show the learned AI models could adapt to real humans. In this work, we have included these approaches as baselines and show that accounting for human beliefs lead to better performance in our setting.
From a technical perspective, our work focuses on understanding and modeling human decision-making. A significant body of literature has explored modeling human behavior. For example, inverse reinforcement learning (IRL)(Ng et al., 2000; Abbeel and Ng, 2004; Ramachandran and Amir, 2007) aims to infer reward functions in Markov decision processes (MDPs) by observing demonstrations of the optimal policy. When the demonstrator is human, these demonstrations may be noisy or exhibit behavioral biases. Several studies(Evans et al., 2016; Shah et al., 2019; Hughes et al., 2020; Zhi-Xuan et al., 2020) have addressed this by incorporating human behavioral biases into the inference process, aiming to infer both the rewards and biases simultaneously. Imitation Learning (Hussein et al., 2017) also seeks to develop models that can mimic human behavior based on demonstrations. Recently, there has been a growing amount of research that incorporates human models into computational and machine learning frameworks (Kleinberg and Oren, 2014; Tang and Ho, 2019, 2021; Masters et al., 2021a, b; Yu and Ho, 2022; Yu et al., 2023; Feng et al., 2024).
From the perspective of modeling human beliefs about others’ intentions, this has been discussed in the literature on level-$k$ reasoning (Stahl II and Wilson, 1994; Gill and Prowse, 2016) in economics and in the theory of mind (Premack and Woodruff, 1978; Westby and Riedl, 2023). For instance, Yu et al. (2006) found that real human behavior aligns closely with level-$1$ and level-$2$ reasoning models in cooperative games. Albrecht and Stone (2018) survey works on autonomous agents modeling the beliefs and intentions of other agents, distinguishing between methods for modeling stationary or changing agent behaviors. These belief models include theory of mind, recursive reasoning models, plan recognition, partially observable Markov decision processes (POMDPs), and others. Baker et al. (2011) propose a Bayesian theory of mind to describe how humans model joint beliefs of state in a partially observable MDP, conducting human experiments in the works of Baker et al. (2009, 2017). Similarly, Wu et al. (2021) apply a human modeling approach in the Overcooked experiment, breaking down the task of delivering a dish into several subtasks, such as picking up an ingredient or placing an ingredient into a pot, where the player infers others’ intentions regarding these subtasks and selects their own accordingly before taking actions. Our belief inference framework differs from previous theory of mind literature by modeling and incorporating human beliefs about AI intentions specifically, using a more realistic human model as the basis for inferring these intentions.
3. Problem Formulation and Methods
In this section, we first formulate the human-AI collaboration framework as a multi-player goal-oriented Markov decision process. We then outline our methods for modeling human behavior and beliefs about AI behavior. Finally, we present our approaches for integrating human behavior and beliefs into the development of collaborative AI agents.
3.1. Decision-Making Environment
We formulate the human-AI cooperative decision-making environment as a multi-agent MDP, represented by $W=\langle S,\alpha,{A_{i}|_{i\in[\alpha]}},R,P\rangle$, where $S$ is a finite set of states, $\alpha$ is a finite set of players, $A_{i}$ is the action set available to player $i$, $P:S\times A_{1}\times\dots\times A_{|\alpha|}\times S\rightarrow[0,1]$ is the transition function that determines the next state given all players’ joint actions, and $R:S\times A_{1}\times\dots\times A_{|\alpha|}\rightarrow\mathbb{R}$ is the joint reward function assigned to all players given their joint actions. Note that in this work we consider the cooperative setting. Therefore, our formulation only incorporates a single reward function $R$ for all players, though this can be easily extended.
While our formulation could address cases with multiple human and AI players, in this work, we focus on a two-player cooperative game, i.e., $\alpha={1,2}$, where one player is a human and the other is the AI. During the decision-making process, neither the human nor the AI knows the other player’s next action or future plans, and they cannot communicate directly. However, they can observe each other’s past actions, allowing them to infer the intentions of the other player and adjust their own actions accordingly.
Agent intentions. In this work, we focus on a setting with a goal-oriented MDP, where there is a set of goals $G={g_{1},\dots,g_{k}}\subseteq S$. The decision-making agent only receives rewards when arriving at one of the goal states, and the goal states are terminal states, i.e., $P(g|g,a)=1\leavevmode\nobreak\ \forall g\in G,a\in A$. Both humans and AI agents can observe past actions but not the specific goals the other agents are trying to achieve. We utilize this formulation to operationalize agent intentions; that is, the goal an agent is trying to reach is defined as the agent’s intention. In the discussion of this paper, we interchangeably use both goals and intentions to refer to intentions.
3.2. Modeling Human Beliefs about AI Intention
Our models are motivated by the level-$k$ framework (Stahl II and Wilson, 1994) in economics. In particular, we start by considering humans as level-00 agents who do not account for others’ behavior in the environment. Unlike previous literature, we address the realistic scenario where level-00 agents may not behave optimally, and we call this model behavioral level-00 agents. To account for human beliefs about AI intentions, [^1] In this work, we sometimes use human beliefs to refer specifically to humans’ beliefs about the intention of the AI agent. we model humans as an extension of level-$1$ agents, who assume other agents are behavioral level-00 agents [^2] Our model can iteratively extend to higher level-$k$ agents; however, we focus on the case with $k\leq 1$. and update their beliefs about AI intentions in a Bayesian manner based on the observation of others’ behavior.
Modeling human behavior. We first describe our models of human behavior under the assumption that humans do not consider other players in the environment (or that they view other players as part of the environment without strategically responding). Our approach to modeling human behavior follows standard methods in the literature and primarily serves as a foundation for developing models of human beliefs. In particular, a human behavior model can be represented as $H:W\rightarrow\Pi$, mapping a given environment $w\in W$ to a policy $\pi=H(w)$. We provide two examples of human behavior models used in our work below.
  • $\bullet$
    Standard model. First consider the standard human behavior model in MDPs, in which the goal of the human is to maximize the expected cumulative reward, and their policy only depends on the current state. The model can be represented by $\pi(a|s)$, indicating the probability of choosing action $a$ at state $s$. For the standard model that assumes decision optimality, humans choose actions maximizing the Q-function, where $Q(s,a)$ indicates the expected cumulative reward if the player takes action $a$ in state $s$ and follows policy $\pi$. $Q(s,a)$ could be calculated by standard reinforcement learning techniques such as value iteration or Q-learning.
  • $\bullet$
    Behavioral level-00 model. We also consider the case where we can learn human behavioral models from their historical behavioral traces using behavioral cloning approaches (Pomerleau, 1988; Torabi et al., 2018). Behavioral cloning is an imitation learning approach that learns a policy from human demonstrations by mapping states to actions using supervised learning methods (Bain and Sammut, 1995). We build a fully connected neural network, where the input is the state encoding and the output is the probability distribution over the action space. The model is trained using the standard gradient descent method with a cross-entropy loss.
Modeling human beliefs. We now describe how we develop models for human beliefs about AI intentions. As summarized earlier, our belief models follows the ideas of level-$k$ reasoning and assume that human decision-makers perceive other agents in the environment as behavioral level-00 agents. In our discussion, we also refer to this belief model as behavioral level-$1$ agents.
More specifically, we model the human belief updating process as a Bayesian inference process, which can be represented as in Equation (1), where $\lambda(g)$ represents the human’s prior belief distributions of AI intentions, and $Pr(s_{t},a_{t}|g)$ denotes the probability of observing $(s_{t},a_{t})$ given the intention (goal) $g$, according to the policy model $\pi(a|s,g)$. This policy model represents how humans perceive the actions of other players. If a human believes the other agent is following the standard model, then the policy is derived from the optimal policy. On the other hand, if humans believe the agent follows the behavioral cloning model, then the policy will be the output of the human model learned from data.
$B(g|(s,a)_{1:t})$ $\propto\lambda(g)Pr((s,a)_{1:t}|g)$ (1)
$=\lambda(g)\prod_{1:t}Pr(s_{i},a_{i}|g)$
$=\lambda(g)\prod_{1:t}\pi(a_{i}|s_{i},g)P(s_{i}|s_{i-1},a_{i-1})$
3.3. Designing Explicable AI Policy
Prior research (Ullman et al., 2009; Baker et al., 2011; Zhi-Xuan et al., 2020) has demonstrated that humans often find it challenging to infer the intentions of other agents based on their behavioral traces, which also aligns with our observations in human-subject experiments (see Section 4 and the appendix for comprehensive results). Therefore, before we delve into the design of collaborative AI, we also explore whether we can leverage the model of human beliefs about AI intentions to design an explicable AI policy, where the AI takes actions that make it easier for humans to infer its intentions.
To train an explicable AI policy, we incorporate the models of human beliefs into the AI training process. Instead of solely aiming to maximize rewards, we also focus on aligning the AI policy to increase the probability that humans can infer AI intentions based on AI actions, leveraging the human belief model to account for the human inference process. More specifically, we utilize self-play to train models. In addition to the original MDP rewards, our approach includes additional bonus rewards when the goal inference based on the human belief model is consistent with the specified goal. Specifically, we set the bonus reward to be proportional to the log likelihood of the belief model inference, $\log(\pi(a|s,g))$. With this setting, the trained AI policy would aim to reach the goal while also maximizing the likelihood of revealing its goals, where the likelihood of the goal inference follows the belief model.
3.4. Designing Collaborative AI Agents
We now describe how we design collaborative AI agents that account for different models of human behavior and beliefs. Later, in the experiments, we illustrate the effectiveness of incorporating models of human beliefs by examining the performance of AI agents paired with humans in human-AI collaboration through simulations and real-world human-subject experiments.
Training methodology. The main idea of our training method is through self-play. We incorporate the models of humans and have AI teammates play with the agents specified by the human models through simulated plays. We use proximal policy optimization (PPO) to train collaborative AI agents.
Collaborative AI agents. Using the methodology above, we have designed several AI agents based on different assumptions of the human counterpart.
  • $\bullet$
    Assuming humans are optimal. We first train an AI agent that learns to collaborate with itself through self-play. This is equivalent to assuming the human counterpart is acting optimally.
  • $\bullet$
    Assuming humans are behavioral level-00 agents. We next train an AI agent that assumes the human is a behavioral level-00 agent, using behavioral cloning to train the behavioral model.
  • $\bullet$
    Incorporating models of human beliefs. Finally we also train an AI agent that incorporates both the models of human behavior and beliefs into the design of AI agents.
In addition to the AI agents trained using our methodologies, we also include two state-of-the-art baselines, Fictitious Co-Play (Strouse et al., 2021) and Maximum Entropy Population-based Training (Zhao et al., 2023), for comparison in our experiments. Both algorithms develop a single best response model in response to a population agents with diverse behaviors, thereby enhancing the robustness and generality of the learned best response model. The details of the setup and implementations are provided in the appendix.
4. Experiments
To evaluate our approaches, we have conducted multiple sets of human-subject experiments through recruiting close to 1,000 participants from Amazon Mechanical Turk (MTurk). The experiments are approved by the IRB of our institution. Workers were paid \$1 base payments with the potential for bonus payments in some experiments. The average hourly rate is approximately \$$\$14$ across all our experiments.
4.1. Experiments Overview
The main purpose of our experiments is to demonstrate that incorporating human beliefs about AI intentions into the design of AI agents can significantly improve the performance of human-AI collaboration. Specifically, we conducted three sets of experiments, beginning with examining the efficacy of our models of human behavior and beliefs, and progressing to evaluate the performance improvements.
In the first experiment, we examine the effectiveness of using behavioral cloning as a model to mimic human behavior. We recruit real-world participants to play against a predefined AI agent to collect their behavioral traces and then evaluate the prediction accuracy of using behavioral cloning to model human behavior. In the second set of experiments, we examine whether we can leverage the model of human beliefs to design an explicable AI policy (as described in Section 3.3). Specifically, we recruit real-world participants, provide them with behavioral traces generated by different AI policies, and evaluate whether the AI policies that account for human belief models indeed produce behavioral traces that make it easier for humans to infer AI intentions. Finally, in the third experiment, we examine whether the collaborative AI that accounts for human beliefs lead to higher performance in human-AI collaboration. Again we recruit real-world human participants, pair them with different AI agents (as described in Section 3.4), and measure the overall team performance.
4.2. Experiment Environment
Our experiments are conducted in a grid world environment with two players and multiple goals. [^3] We conducted experiments in two variations of the environment. In the first variation, we designed the environments such that the two players are not in the same grid world. However, each player has access to full information about the environment and the actions of the other player. This variation abstracts away the interdependency of agent actions, meaning the agents’ actions do not directly influence each other. This allows us to focus on how humans and AI reason about each other’s goals and intentions. In the second variation, we consider a more complex and realistic situation where both players are in the same grid world environment and may collide with each other. We obtained qualitatively similar results in both variations. Due to space constraints, we have included the setup and results of the second variation in the main paper and moved the results of the first variation to the appendix. More specifically, as shown in Figure 2, the grid world is 8 by 8 in size and contains the positions of both players and four possible goals. The players can choose to move {Up, Down, Right, Left} or stay in their current grid. They can see each other’s positions and take actions simultaneously. Each player can navigate to two of the four goals. In particular, the human player can only reach one of the two goals colored blue, and the AI player can only reach one of the goals colored green. When both agents reach the same type of goal (e.g., both reaching a ”star” or both reaching a ”triangle”), they earn positive points. However, they do not earn points if they reach different types of goals or if they collide with each other (move into the same position). We set the maximum number of actions to be 20.
Figure 2. The decision-making environment in simulations and human-subject experiments. The human-AI team will receive a reward by reaching the same type of goals (star or triangle) within the time limit. They will receive no reward if they collide with each other or reach different types of goals.
4.3. Experiment 1: Evaluating Behavior Models
In our first experiment, our goal is to examine the effectiveness of utilizing behavioral cloning to model human behavior. Note that his approach has been proposed and examined in various domains Carroll et al. (2019). Our main purpose to conduct this experiment is to ensure whether this method works well in our setting, as our proposed approaches in developing human belief models and collaborative AI are built on top of models of human behavior.
Experiment setup. We recruited 190 workers from Amazon Mechanical Turk. Each recruited worker was asked to play 30 navigation games within a grid world with a predefined AI model, as shown in Figure 2. To collect diverse data, we randomly paired recruited workers with one of three AI models: a random AI (selecting a random goal and taking the shortest path), a self-play AI (trained using self-play), and an optimal AI (optimized via joint optimization).
In this experiment, our goal is to develop a model of human behavior that outputs a sequence of actions given a goal. Therefore, in Experiment 1, we provided participants with additional instructions about which goal they should reach in each round, and the AI models controlled the other player to move towards the same type of goals. In our later Experiment 3, we did not provide these instructions, and humans needed to infer the goal of the AI agent.
Experiment results. We divided the collected data of human actions into three sets: training, validation, and testing. The training set comprised data from 152 workers, including approximately 136,000 instances of user decisions, while the validation and testing sets each contained data from 19 workers, amounting to around 17,000 instances of user decisions each. We employed a 4-layer Multilayer Perceptron (MLP) model, where the input is the current environment layout, and the output predicts the next human action. We fine-tuned the hyperparameters, such as learning rate, hidden layer size, and L2 penalty, based on validation errors.
Table 1. The prediction accuracy for human behavior was evaluated assuming optimal behavior and using a data-driven model in Experiment 1. The results show that human behavior deviates significantly from optimality, and our data-driven model provides a much more accurate prediction of human behavior.
Training Accuracy Validation Accuracy Testing Accuracy
Assuming Optimal Behavior 0.4498 0.4327 0.4459
Data-Driven Model 0.8547 0.7831 0.7899
Our results suggest that human behavior is significantly away from being optimal. Even when human participants are provided with suggested goals during data collection, there is only a $55.0\%$ chance that both players will reach the same type of goals across all treatments. Besides, there is a notable chance of two players colliding (about $15.8\%$) or ending in different types of goals (about $25.6\%$). We compared the performance of our learned model to a model predicated on optimal agent behavior, defined as taking the shortest path to the goal. The training, validation, and test accuracies of both models are presented in Table 1. These results clearly reveal that human behavior deviates significantly from the assumed optimality. This deviation highlights the importance of incorporating a realistic model of human behavior into human-AI cooperation frameworks.
4.4. Experiment 2: Evaluating Belief Models
We next evaluate our model of human beliefs. In particular, we examine whether an AI policy that incorporates human behavior models can produce actions that are more explicable to humans, i.e., easier for humans to infer AI intentions based on AI actions. [^4] We have also directly tested the accuracy of our belief model in inferring human beliefs by comparing its predictions with direct solicitations of users’ inferences about others’ behavior. The results, included in Appendix A.3, suggest that our model is more accurate in predicting human beliefs compared to the baselines. However, they also demonstrate that human beliefs are generally quite noisy—if we randomly draw a behavior trace from another agent and ask humans to predict the goal, human prediction accuracy is close to random guessing. Therefore, we shift our focus in this experiment to whether AI agents can design their action plans to make it easier for humans to infer the goal.
Experiment setup. In this experiment, each recruited participant is presented with the behavioral traces from both agents in the environment, as shown in Figure 3. Each behavioral trace is a sequence of actions generated by an AI policy for a player. Participants are then asked to infer the intention of each agent, i.e., the goal each agent is trying to reach. If participants can better infer the goal from the behavior generated by a certain AI policy, it indicates that the AI policy is more explicable in that it generates actions that make it easier for humans to infer AI intentions.
Figure 3. The experiment interface in Experiment 2. Participants are given the behavioral traces of the agents and are asked to infer the goal each agent aims to reach.
AI policy. We now describe the set of AI policies that are used for the evaluations. As baselines, the first two AI policies are not intended to be explicable.
  • $\bullet$
    Inexplicable-Optimal-L0: This AI policy that takes the shortest path towards the goal.
  • $\bullet$
    Inexplicable-Behavioral-L0: This AI policy takes actions that follow the data-driven behavioral model learned in experiment 1.
We then include two AI policies that take actions intended to be explicable. Note that to design AI policies to be explicable, we need to account for human belief models in the design of AI policies, as the AI will try to take actions that make it easier for humans to infer the goal, assuming models of human beliefs that describe how humans make inferences. The two belief models we account for are: (1) the classical level-$1$ model, where the human assumes the other agent is taking optimal level-00 actions, and (2) the behavioral level-$1$ model, where the human assumes the other agent is taking behavioral level-00 actions as described in Section 3.2. Coupled with these two belief models, and using the methodology described in Section 3.3, we can train the following two explicable AI policies.
  • $\bullet$
    Explicable-Optimal-L1: This AI policy is trained to be explicable but assumes that humans belief model is the classical level-$1$ model, where they assume the other agent is optimal.
  • $\bullet$
    Explicable-Behavioral-L1: This AI policy is trained to be explicable but assumes that human belief model is the behavioral level-$1$ model, where they assume the other agent is behavioral level-00.
Figure 4. The results of the human-subject experiments in Experiment 2 show that when using our approach to design explicable AI, the accuracy for humans to infer AI intentions is higher than with other AI policies. Moreover, when coupled with our belief model, it results in an AI policy that makes it the easiest for humans to infer AI intentions.
Experiment results. We recruited 300 workers from Amazon Mechanical Turk, randomly assigning them to one of four treatments, where each treatment represents showing participants traces generated by one of the AI policies above. Each participant was tasked with identifying the goals of the player in 30 different scenarios. The length of actions is drawn from the range 4 to 8. To incentivize effort, participants were awarded a \$$\$0.03$ bonus for each correctly identified goal.
The experiment results are shown in Figure 4. First, we observe that when the AI is not attempting to be explicable, humans generally struggle to infer AI intentions, with the accuracy of correctly inferring the goal for both inexplicable policies being around 55% to 60%, only slightly better than random guessing. However, if we use our methodology from Section 3.3 to train the AI policy to be explicable, it significantly improves the accuracy of humans in inferring AI intentions. Moreover, when the training is coupled with our belief models, the accuracy for humans to infer AI intentions is the highest, exceeding 70%.
4.5. Experiment 3: Evaluating AI Design
We now examine our design of collaborative AI agents. Our goal is to demonstrate that the collaborative AI design that accounts for human beliefs could lead to improved human-AI collaborative performance. We utilize both simulations and human-subject experiments to examine the effectiveness of our design.
Table 2. Simulation results of collaborative performance over 10,000 testing cases in Experiment 3. The column players represent different AI agents, and the row players represent different simulated human models. The results demonstrate that while the zero-shot methods (FCP and MEP) show consistent and stable performance, when human agents possess belief models and infer AI intentions to act accordingly, designing AI that accounts for human beliefs brings significant performance improvements.
Human Model AI Agent
Self-Play FCP MEP Behavior-AI Belief-AI
Self-play 0.6245 0.6136 0.6131 0.5250 0.6177
Behavior Model 0.4902 0.5782 0.5927 0.6334 0.5755
Behavior&Belief 0.6411 0.6255 0.6484 0.6574 0.7675
Experiment setup. The setup of our Experiment 3 is the same as Experiment 1, shown in Figure 2. But human participants will not receive any hint about which goal their teammate or themselves are suggested to reach.
Collaborative AI design. As described in Section 3.4, we train collaborative AI agents using three different corresponding human models.
  • $\bullet$
    Self-Play: the collaborative AI agent that is trained assuming they are playing with itself.
  • $\bullet$
    Behavior-AI: the collaborative AI agent that is trained assuming the human partner is the behavioral level-00 agent.
  • $\bullet$
    Belief-AI: the collaborative AI agent that is trained assuming the human partner is the behavioral level-$1$ agent, i.e., it accounts for human belief models into the design of collaborative AI.
In addition to the three collaborative AI agents trained using our methodology, we also include two state-of-the-art baselines for comparison. These methods leverage zero-shot population-based training (PBT) techniques. While they do not explicitly account for specific models of human behavior, they are designed to be robust in collaborating with different teammate behaviors.
  • $\bullet$
    FCP: Fictitious Co-Play is a variant of fictitious play proposed by Strouse et al. (2021). It consists of two stages: in the first stage, a population of agents is trained independently using different random seeds and checkpoints; in the second stage, a model is trained as the best response to the agent population. In our experiment, we use the model trained with 8 partners.
  • $\bullet$
    MEP: Maximum Entropy Population-based training is proposed by  Zhao et al. (2023). The agent population is trained using population entropy bonus to increase the diversity of agents, then a single best response model is trained with respect to the agent population. In our experiment, we use the model trained with 8 partners.
Simulations. We first run simulations to evaluate the performance of different pairings of human models and collaborative AI designs, where we assume the human agent is optimal (self-play), an behavioral level-00 agent (Behavioral Model), and behavioral level-$1$ agent (Behavior&Belief). The evaluation is based on $10,000$ randomly generated environments which filter out cases where the distance between the two goals for the same player is smaller than 3 to allow models to adjust their behavior based on the inference of their teammate’s actions.
We partner the collaborative AI agents with the three simulated human agents and measure the collaborative performance. The simulation results are shown in Table 2. For both the FCP and MEP baselines, despite the absence of human data in their training processes, they outperform the self-play AI when paired with the human behavior model, demonstrating their adaptability to working with teammates exhibiting unknown behavior. However, we also observe that incorporating the appropriate human behavior model directly into the training process can further enhance human-AI collaborative performance, as demonstrated by the superior performance of Behavior-AI when paired with the behavior model. Finally, when working with a behavioral level-1 agent that attempts to infer AI intentions, our proposed Belief-AI stands out, significantly outperforming other AI agents, which perform similarly to each other.
These results highlight the importance of selecting appropriate human models for training collaborative AI. Furthermore, under conditions where humans attempt to infer AI intentions and act accordingly, designing AI that accounts for human beliefs can significantly improve human-AI collaborative performance.
Human-subject experiments. While the simulation results are promising, the significant performance improvement from incorporating human beliefs about AI intentions into AI design is based on the assumption that humans do indeed infer AI intentions and act accordingly. To examine whether this performance improvement carries over to collaboration with real humans, we recruited 400 workers and randomly assigned them into five treatment groups, each interacting with a different AI design. Participants were tasked with playing 30 games, preceded by three tutorial games designed to familiarize them with the gameplay. Participants could receive a \$0.05 bonus payment for reaching the same type of goal as the AI agent in each of the 30 tasks.
The experiment results are shown in Figure 5. They demonstrate that the collaborative AI trained with models of human behavior and beliefs achieved the highest collaborative performance when working with real humans. This outcome not only validates our AI design but also suggests that real human actions align with the behavior and belief models in this environment. Statistical analysis shows significant differences in the performance of AI models paired with humans, with $p<0.0001$ for comparisons between self-play AI and AI accounting for human behavior, and $p<0.0001$ for comparisons between AI accounting for human behavior versus accounting for both behavior and beliefs. The differences between MEP and FCP, as well as between FCP and behavioral-AI, are not significant ($p>0.1$). The difference between self-play AI and any other treatment is significant ($p<0.0001$). These results demonstrate that incorporating human beliefs into the design of AI agents enhances collaborative performance when working with real humans.
Figure 5. The human-AI collaborative performance in our human-subject experiment in Experiment 3 shows that designing collaborative AI that accounts for human beliefs (Belief-AI treatment) significantly outperforms all other baselines when working with real humans.
5. Conclusion
Our work explores the impact of incorporating human beliefs about AI intentions into the design of collaborative AI, aiming to improve collaboration between humans and AI. By developing models that account for human beliefs regarding AI intentions and integrating these models into AI design, we have observed improved performance in human-AI collaboration. Our approach, validated through human-subject experiments, demonstrates that AI agents designed with an understanding of human behavior and beliefs over AI intention can be more effective in working with humans. This suggests a potential path forward for creating AI systems that are better aligned with human partners, making collaborative tasks more efficient and effective.
Generalization and limitations. Our approach has demonstrated effectiveness through human-subject experiments conducted in commonly used grid world environments of varying complexity. However, like most human-subject studies, our findings are limited to the chosen environments and conditions. To enhance the external validity of our approach, it is important to evaluate its performance across different domains and more realistic settings beyond grid worlds, such as dynamic, real-world applications involving physical robots or complex multi-agent systems.
Additionally, our approach leverages historical data to construct models of human behavior and beliefs, which implicitly assumes that human behavior and beliefs remain stable over time. This may not hold true in real-world scenarios where human behavior can change in response to prolonged exposure to AI or due to evolving task contexts. Future research could focus on developing adaptive models that can learn and update human behavior and belief patterns in real-time to address this limitation. Investigating the scalability of these models to larger, more diverse datasets and environments is also important and interesting for assessing the generalizability and robustness of our approach.
Societal impacts. Our work underscores the importance of integrating human behavior and beliefs about AI into the design of AI systems to enhance human-AI collaboration. As AI systems become increasingly autonomous and capable, it is critical that these systems are designed with an understanding of human behavior and beliefs. This alignment can help improve trust, reduce errors in collaborative settings, and foster more effective interactions.
However, there are also potential ethical considerations and potential risks due to the dual-use of our approaches. In particular, the same methods used to improve human-AI collaboration could be misused, such as designing AI systems that manipulate or deceive users by exploiting their expectations or beliefs. This could undermine trust and create systems that harm user welfare. To mitigate these risks, one potential approach is to develop and enforce ethical guidelines that prioritize transparency in AI models. Providing clear and understandable explanations for AI behavior can help users build appropriate levels of trust and reliance. Moreover, promoting responsible AI design principles that incorporate user consent, control, and oversight are also important in ensuring that the technology serves societal benefits rather than causing harm.
{acks}
This work is supported in part by a J.P. Morgan Faculty Research Award and a Global Incubator Seed Grant from McDonnell International Scholars Academy.
Appendix A Additional Experiments
Besides experiments in Section 4, we also conduct additional experiments in a simple grid world environment with multiple goals to illustrate our study of humans’ beliefs about AI behavior. We also explore how to improve the model accuracy of behavioral cloning approach.
A.1. Experiment Environment: Grid World with Multiple Goals
We use grid worlds of size of $6\times 6$ in both simulations and human experiments. Similar to the environment setup in Section 4.2, the grid world contains a start position, two goal positions, and some blocked positions that the player cannot enter. The player needs to move from the start position towards one of the goal positions. The player can choose to move {Up, Down, Right, Left}. The player will get a positive reward upon reaching the goal, and we set the maximum number of actions to be 20.
We first recruit participants to engage in a single-player game and record their behavior. We then build human behavior model using behavioral cloning. To evaluate our proposed approaches in modeling human beliefs and designing cooperative AI, we conducted additional two sets of experiments. In the second set of experiments, we provide participants with different traces of actions from other agents, again in a single-player game, and ask participants to infer the goal of the agents. This experiment helps us evaluate whether our belief model leads to better predictions of human beliefs over others’ behavior. Afterwards, we conduct a two-player game in the third experiment, where humans are paired with different AI agents to examine the team performance of our design of collaborative AI.
The interfaces of our experiments can be seen in Figure 6a, 6b, and 6c. The detailed descriptions of the experiment and interfaces are included in Appendix C.
(a) Experiment 4.
(b) Experiment 5.
(c) Experiment 6.
Figure 6. Human-subject experiment interfaces. In Experiment 4, each participant is asked to control the player to move to the goal (red star). In Experiment 5, each participant is provided a trace of the behavior by another agent, and is asked to infer which goal the agent is trying to reach. In Experiment 6, each participant is playing with an AI agent in separate environments. The participants only receive bonus rewards by reaching the same goal (red star or green triangle) as the AI agent.
A.2. Experiment 4: Evaluating Behavior Models in Single-Player MDP
Similar to Experiment 1, we recruited 200 workers from Amazon Mechanical Turk. Each recruited worker was asked to play 15 navigation games within a grid world, as shown in Figure 6a. Our goal is to leverage the collected data to create a data-driven model of human behavior. We divided the collected data of human actions into three sets: training, validation, and testing. The training set comprised data from 160 workers, including approximately 70,000 instances of user decisions, while the validation and testing sets each contained data from 20 workers, amounting to around 8,800 instances of user decisions each. The model and parameter tuning were similar to what we did in Experiment 1.
Evaluation of data-driven behavior models. The training, validation, and test accuracies of assuming optimal behavior and behavioral models are presented in Table 3. Results show that data-driven model could predict human behavior more accurately than assuming humans are optimal.
Table 3. The prediction accuracy for human behavior for different human models in Experiment 4.
Training Accuracy Validation Accuracy Testing Accuracy
Assuming Optimal Behavior 0.7266 0.6964 0.7131
Data-Driven Model 0.9189 0.8136 0.8422
A.3. Experiment 5:
Evaluating Human Belief Models
We developed human belief models (behavioral level-1) similar to previous experiments. We examine whether belief inference using behavioral model aligns with real humans, and whether we can design AI behavior such that it is easier for humans to infer the goal of the AI agent.
Experiment setup. The experiment setup is presented in Figure 6b. The grid world contains a starting position and two goal positions. For each participant, we show them a trace of behavior from another agent and ask the participant to infer which of the two goal the agent is trying to reach.
Experiment 5.1: Examining the belief models. We recruited 200 workers from Amazon Mechanical Turk to compare standard level-$1$ and behavioral level-$1$ model. Each worker was asked to infer the goal for 25 behavioral traces. We compared the performance of two belief models and the worker accuracy, as shown in Table 4. As we can see from the table, humans are generally poor at inferring the goal of other agents: their accuracy only reaches $59.37\%$ in inferring the goal of the other agent. Moreover, the behavioral level-$1$ model, which accounts for the human behavior model in the belief model, captures human beliefs better than the standard level-$1$ model, which assumes human behavior is optimal.
Table 4. Performance comparison between Bayesian inference framework assuming standard model and human behavior model.
Consistency to Human Predictions Cross Entropy Loss
Standard level-$1$ 0.4977 0.9284
Behavioral level-$1$ 0.5764 0.7506
True goals 0.5937 0.6631
Figure 7. Human evaluations of belief inference accuracy regrading AI goals.
Experiment 5.2: Developing explicable AI policy. As demonstrated in Experiment 5.1, humans generally struggle to infer the goals of other agents based on others’ behavior. To design AI agents which makes inference of their goals easier by humans, we train AI agents to maximize the likelihood that humans can accurately infer these goals from their behavior.
We recruited 400 workers from Amazon Mechanical Turk, randomly assigning them to one of two treatments, to assess the behavior of AI agents. Each participant was tasked with identifying the goals of the player in 30 different scenarios. Participants were awarded a \$$\$0.03$ bonus for each correctly identified goal. Figure 7 displays the results of human evaluation. The findings indicate that humans achieve higher accuracy in inferring the correct goals of AI agents when employing Behavioral level-$1$ model.
A.4. Experiment 6:
Evaluating Collaborative AI Design
Utilizing developed models of human behavior and beliefs, we follow the same methodology in Experiment 3: train different collaborative AI, pair them with different human models, and examine the collaborative performance in simulations and human subject experiments (via recruiting 300 workers).
The setup of our Experiment 6 is shown in Figure 6c. The goal for the human-AI team is for both agents to reach the same goal in their own environments (both reaching “red star” or both reaching “green triangle”) within a time limit. The team will not get points if they reach different goals or one of the players fails to reach any goal.
Simulations results are shown in Table 5. Simulations indicate that collaborative performance is highest when an AI agent is paired with the human model used to train it. Figure 8 presents the average collaborative reward in human-subject experiments. Statistical analysis revealed significant differences in the performance of AI models when paired with human participants, with $p$-values of $0.0166$ for comparisons between the treatments self-play and Behavior-AI, and $p<0.0001$ for Behavior-AI versus Belief-AI. These results demonstrate that incorporating human beliefs into the deign of AI agents enhances collaborative performance when working with real humans.
Table 5. Simulation results of human-AI collaborative rewards in Experiment 6. Columns players are different AI agents, and row players are different simulated human models.
Human Model AI Agent
Self-play Behavior-AI Belief-AI
Self-play AI 0.7828 0.4780 0.6164
Behavior 0.5926 0.7584 0.4552
Behavior&Belief 0.6919 0.7268 0.7813
Figure 8. Average collaborative reward in human-subject experiment 6.
A.5. Extensive Results of Behavioral Cloning
We investigate techniques to enhance the accuracy of behavioral cloning models. Our experiment results reveal that human actions depend significantly on historical actions. We explicitly include the action history of both AI and human players as part of the model input to determine if this improves the prediction of human actions. The findings, presented in Figure 9, indicate that incorporating action history enhances the accuracy of our behavioral cloning model predictions. Additionally, our belief model consistently boosts prediction accuracy across nearly all scenarios.
(a) Behavioral cloning models.
(b) Behavioral cloning & belief models.
Figure 9. Evaluation of model accuracy on human-subject experiment data through incorporation of action history.
Appendix B Implementation Details
All our experiments are conducted on a cluster of 40 CPU cores (Intel Xeon Gold 6148 CPU @@@@ 2.40GHz), 2 GPUs (NVIDIA Tesla V100 SXM2 32GB), and a maximal memory of 80GB. For completeness, we provide the details of the implementation and the tuning of hyperparameters below.
State encoding of grid world. For a grid world of size $N\times N$ and contains $k$ different objects, we use $k$ channels to encode the environment state. Each channel is a binary matrix of size $N\times N$, indicating whether it contains specific objects. For example, in a single-player environment with one goal (environment in Experiment 4), we encode space, blocks, the start position, and goal positions as four channels. Thus the input state is of size $4\times 6\times 6$ before flattening. The state encoding is used as input of our behavioral cloning models and AI agents.
Behavioral cloning for grid world of size $6\times 6$. We construct a neural network with four fully connected layers to model human behavior. The input is the state encoding, represented as a matrix of dimensions $4\times 6\times 6$, and the output is the probability of taking various actions. The network parameters are initialized using the Glorot uniform initializer. We employ the Adam optimizer and cross-entropy loss to optimize the neural network parameters. The batch size is set to $1024$, and the maximum number of training epochs is set to $1000$. Training is halted early if the validation loss begins to increase. The hyperparameters are adjusted based on the performance on the validation dataset. The number of nodes for each hidden layer is tuned within the range $\{64,128,256,512,1024\}$, the initial learning rate within the range $\{0.001,0.003,0.01,0.03,0.1\}$, and the L2 penalty within the range $\{0,0.01,0.03,0.1\}$. Each training process takes $10$ to $30$ minutes depends on hyperparameters. In our reported results, we choose hidden layer size of $512$, learning rate $0.01$ and L2 penalty $0.01$.
Behavioral cloning for grid world of size $8\times 8$. The neural network structure mirrors that used for the grid world of size $6\times 6$. The input is the state encoding, represented as a matrix with dimensions $8\times 8\times 8$, and the output is the probability of taking various actions. The initialization, optimizer, loss function, batch size and training process are consistent with those used in the grid world of size $6\times 6$. The maximum of training epochs is set to $3000$. The hyperparameters are adjusted based on the performance on the validation dataset. The number of nodes for each hidden layer is tuned within the range $\{1024,2048,4096\}$, the initial learning rate within the range $\{0.001,0.003,0.01,0.03,0.1,0.3\}$, and the L2 penalty within the range $\{0,0.01,0.1\}$. In our reported results, we choose hidden layer size of $2048$, learning rate $0.01$ and L2 penalty $0.01$.
AI agent training for grid world of size $6\times 6$. We utilize the Proximal Policy Optimization (PPO) and actor-critic framework for our AI model. Both the actor and critic networks employ four fully connected layers, maintained at the same size. The input is the state encoding of a two-player game, which includes two game states and the index vector (the input vector length is $290$). The critic network outputs a single value, while the actor network produces a distribution over the action space. The batch size is set to $256$, and the maximum number of training episodes is limited to $1000$. The discount factor is set to $0.99$. For PPO optimization, the clipping parameter $\epsilon$ is fixed at $0.2$, and the number of PPO epochs is $10$. The number of nodes for each hidden layer is tuned within the range $\{64,128,256,512,1024\}$, and the initial learning rate is adjusted within the range of $\{0.001,0.003,0.01,0.03,0.1\}$. Each training process takes about $2$ to $6$ hours depends on hyperparameters. In comparisons between the self-play AI and the AI trained to play with a human model, we use the same model size (hidden layer size of $256$) and the same initial learning rate of $0.01$.
AI agent training for grid world of size $8\times 8$. The model structure mirrors that of the grid world of size $6\times 6$. The input size is adjusted to accommodate the new state representation (the input vector length is $514$). The training batch size, discounting factor, maximum number of training epochs and clipping parameters remain unchanged. The number of nodes for each hidden layer is tuned within the range $\{256,512,1024,2048\}$, and the initial learning rate is adjusted within the range of $\{0.01,0.03,0.1,0.3,0.5\}$. In comparisons between the self-play AI and the AI trained to play with a human model, we consistently use the same model size (hidden layer size of $512$) and the same initial learning rate of $0.1$.
Selection of population size.
For FCP and MEP algorithms used in Section 4, the population size will directly influence the model performance. We vary the number of agents in our simulations as shown in Figure 10. The model trained with 8 partners is chosen for use in further simulations and human-subject experiments for comparative analysis.
Figure 10. Average team reward for models trained across different population sizes using the FCP and MEP algorithms.
Appendix C Human Experiment Details
We include more details about human-subject experiments. The experiment interface is shown in Figure 11. We conduct four human-subject experiments, and recruited 1990 workers from Amazon Mechanical Turk in total. Table 6 lists the number of workers, the number of tasks per worker and the number of treatments in each experiment. Table 7 contains the demographic information of all the workers.
(a) Experiment 1 & 3 interface: humans play games with AI agents in grid worlds of size $8\times 8$.
(b) Experiment 2 interface: humans infer goals of other players in grid worlds of size $8\times 8$.
(c) Experiment 4 interface: humans play single-player navigation games in grid worlds of size $6\times 6$.
(d) Experiment 5 interface: humans infer goals of other players in grid worlds of size $6\times 6$.
(e) Experiment 6 interface: humans play games with AI agents in grid worlds of size $6\times 6$.
Figure 11. Human-subject experiment interfaces.
Table 6. Number of workers in each human-subject experiment.
Workers Tasks per Worker Treatments Base Payment Bonus per Task
Experiment 1 190 30 3 \$1.00 \$0.05
Experiment 2 300 30 4 \$1.00 \$0.03
Experiment 3 400 30 5 \$1.00 \$0.05
Experiment 4 200 15 1 \$1.00 0
Experiment 5.1 200 25 1 \$1.00 0
Experiment 5.2 400 30 2 \$1.00 \$0.03
Experiment 6 300 20 3 \$1.50 \$0.05
Table 7. Demographic information of all the participants in our experiments.
Group Category Number
Age 20 to 29 714
30 to 39 974
40 to 49 176
50 or older 72
Other 54
Gender Female 639
Male 1319
Other 32
Race / Ethnicity Caucasian 1772
Black or African-American 105
American Indian/Alaskan Native 36
Asian or Asian-American 29
Spanish/Hispanic 11
Other 37
Education High school degree 66
Some college credit, no degree 22
Associate’s degree 36
Bachelor’s degree 1457
Graduate’s degree 353
Other 56